For an ideal I of a commutative ring R, define

√I {x belongs to R |x^n belongs to I |for some n ∈N}. Show that

i) √I is an ideal of R.

ii) I ⊆ √I.

iii) I ≠ √I in some cases.

b) Is R/I×R/J=R×R/I×J ,for any two ideals I and J of a ring R ? Give reasons for your

answer


1
Expert's answer
2021-05-07T09:15:08-0400

1) Let "x, y\in\sqrt{I}" . There exist "n,m\in \mathbb{N}" such that "x^n,y^m\in I". Then, "(x+y)^{n+m}=\sum\limits_{k=0}^{n+m}\binom{n+m}{k}x^ky^{n+m-k}=y^m\sum\limits_{k=0}^{n}\binom{n+m}{k}x^ky^{n-k}+x^n\sum\limits_{k=n+1}^{n+m}\binom{n+m}{k}x^{k-n}y^{n+m-k}\in I"

Hence "(x+y)\in \sqrt{I}"

For any "a\in R\ (ax)^n=a^nx^n\in I" , hence "ax\in \sqrt{I}"

Therefore, "\sqrt{I}" is an ideal.


2) "\forall x\in I\,\forall n\in\mathbb{N}x\, x^n\in I" , hence "x\in \sqrt{I}" and "I\subset \sqrt{I}" .


3) Let "R=\mathbb{Z}", "I=(4)", "\sqrt{I}=\{x\in \mathbb{Z}: \exists n\in\mathbb{N}\ 4|x^n\}=\{x\in \mathbb{Z}: 2|x\}=(2)". So, "I\ne\sqrt{I}" .


4) "R/I\times R/J=(R\times R)/(I\times J)"

Indeed, "I\times J" is an ideal in "R\times R":

"\forall i_1,i_2\in I \forall j_1,j_2\in J\ i_1+i_2\in I, j_1+j_2\in J", hence "(i_1+i_2,j_1+j_2)\in I\times J"

"\forall (i,j)\in I\times J\ \forall (a,b)\in R\times R\ ai\in I,\ bj\in J,\ (ai,bj)\in I\times J"


Define "f:R/I\times R/J\to(R\times R)/(I\times J)" by the formula "f(x+I,y+J)=(x,y)+(I\times J)"


This definition is correct, since if "x_1-x_2=i\in I" and "y_1-y_2=j\in J" then

"f(x_1,y_1)=f(x_2+i,y_2+j)=(x_2+i,y_2+j)+(I\times J)=(x_2,y_2)+(I\times J)=f(x_2,y_2)"

Since "I,J" are ideals in "R" , and "I\times J" is an ideal in "R\times R" , we have

"f(x_1+x_2+I,y_1+y_2+J)=(x_1,y_1)+(x_2,y_2)+(I\times J)=f(x_1+I,y_1+J) + f(x_2+I,y_2+J)"

"f(ax+I,by+J)=(ax, by)+(I\times J)=(a,b)f(x,y)"

therefore, f is homomorphism.


f is surjective, since for any class "(x,y)+I\times J\in R\times R/(I\times J)" we have "(x,y)+I\times J=f(x+I,y+J)"


f is injective, since if "f(x+I,y+J)=(0,0)+(I\times J)", then "x\in I" and "y\in J" and "(x+I,y+J)=(0+I,0+J)" is zero in "f:R/I\times R/J" .

Therefore, f is an isomorphism.